OpenStudy (anonymous):

Given the linear equation y = 2x – 5, find the y-coordinates of the points (-2, ), (4, ), and (-1, ). Please show all of your work

OpenStudy (niel):

simple but effective way to do it: place -2 where the x is and do the math. you will find y if x is -2 .... now do the same for the others and yull be good

OpenStudy (anonymous):

for y coordinates set x =-2, 3, -1 for x = -2: y = 2(-2)-5 = -9 (-2,-9) for x = 4: y = 2(4)-5 = 3 (4,3) for x =-1: y = 2(-1)-5 = -7 (-1,-7)

OpenStudy (niel):

give a man , or woman a fish, feed him/her for a day....

OpenStudy (anonymous):

1) y=2x-5 y=2(-2) - 5 y=-4-5 y= -9 2)y=2x-5 y=2(4)-5 y=8-5 y=3 3)y=2x-5 y=2(-1)-5 y=-2-5 y=-7

OpenStudy (niel):

great job and i think you did it on your own... right?lol

OpenStudy (anonymous):

thanks how would i plot this

OpenStudy (niel):

y = 2x -5 y = mx+b m = 2 in this case which is the slope b = -5 which is where the line hits the y axis so you have starting at point 0,2 mark it then go up 2 and to the right 1

OpenStudy (niel):

its the same method for all three